# Exploring Large-Volume GAGG Scintillators for use in MeV Gamma-Ray Astrophysics

### v1

@@ Abstract
Gamma-ray astrophysics in the MeV band is an exciting field in astronomy due to its potential for multi-messenger astrophysics. It has, however, remained under-explored when compared to other wavelengths. One reason for this observational gap is the difficulties with measuring these high-energy photons and the requirement of large amounts of detection material. In this work, we investigate the usage of large-volume GAGG scintillators for use as a calorimeter in future MeV telescopes. We developed a $5\times 5$ array calorimeter utilizing $1\times 1\times 6\ \mathrm{cm}^{3}$ GAGG crystals with C-series SiPM readout. The calorimeter was tested at the High Intensity Gamma-ray Facility (HIGS) with monoenergetic beams ranging from $2-25\ \mathrm{MeV}$. Finally, we also investigated larger $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals and characterized their response across their depth when their surface treatment is either polished or frosted.
@@ 1 Introduction
The gamma-ray astrophysical sky in the MeV range remains largely under-explored. The most recent large-scale mission dedicated to measuring and imaging the gamma-ray sky was the Compton Gamma-Ray Observatory (CGRO), which was de-orbited in 2000 [1]. Since then, the community has proposed several Compton and/or pair-production instruments to address what is known as the ‘MeV gap’. Most pair-production instruments are based on a two-subsystem design: a tracker to trace out the movement of the cascade of particles from the interaction of the gamma ray and a calorimeter to capture the ensuing electromagnetic showers. This tracker-calorimeter design is utilized by current missions such as the Fermi Space Telescope-LAT [2] and AGILE [3]. Missions that are in the concept and development phase include AMEGO-X [4], ASTROGAM [5], SMILE-2+ [6], VLAST [7], and GECCO [8].
@@
GAGG CsI:Tl
Light Yield (ph/keV) 45 54
@@
Density (g/$\mathrm{cm}^{3}$) 6.3 4.51
Decay Time (ns) $\sim$140 1000
@@ 2 Development of a $5\times 5$ array GAGG Calorimeter
In this work, we utilize high light-yield GAGG. The prototype in this section uses $1\times 1\times 6\ \mathrm{cm}^{3}$ crystals, each one wrap with a diffuse white reflector (Tetratex) [19], and place a $6\times 6\ \mathrm{mm}^{2}$ C-series SiPM by onsemi [20] on each end of the crystal. Fig. 1 shows the bare crystal and a fully developed crystal element.
@@ 3 High-Energy Gamma-Ray Response of the Calorimeter
In April 2022, the calorimeter was taken to the Triangle Universities Nuclear Laboratory (TUNL) High Intensity Gamma-Ray Source (HIGS) facility. The facility can produce a monoenergetic gamma-ray beam with a diameter of $0.5\ \mathrm{cm}$. The experiment utilized beam energies of 2, 5, 7, 15, and 25 MeV. Fig. 5 shows the experimental setup. Since this experiment was conducted as an add-on to the AMEGO/ComPair campaign [23], the GAGG is placed downstream of the ComPair instrument. The ComPair instrument consists of 3 subsystems: 6 layers of $0.5\ \mathrm{mm}$ of Si, $2\ \mathrm{cm}$ of CZT, and $8.35\ \mathrm{cm}$ of CsI.
@@
During the experiment, we placed the calorimeter in different positions and angles to the beam. This is to change the interaction profile of the gammas with the calorimeter. Fig. 6 shows the summed response to different irradiation configurations of a 5 MeV beam. Side irradiation refers to when the beam is shot into the center column of the calorimeter while front irradiation is where the beam is aimed into the $1\times 1\ \mathrm{cm^{2}}$ face of the center crystal in the array. We note that the difference in the full calorimeter response is relatively invariant in that the ratio between the escape peaks, continuum, and full-energy peak are very similar. This is likely due to the near-cubic geometry of the calorimeter which results in a uniform sensitivity to source location.
@@
Figure 7: Crystal-to-crystal response to 5 MeV beam while the calorimeter is oriented at different poses.
Fig. 8 plots all the acquired spectra from the HIGS campaign. Full energy peaks are clearly visible for the 2, 5, and 7 $\mathrm{MeV}$ beams. Also featured are the single and double escape peaks. However, runs with 15 and 25 $\mathrm{MeV}$ do not present any evidence of escape peak, likely due to the low full-energy efficiency and full-containment of the ensuing spallation.
@@ 4.1 Collimated Depth Measurements
In this section, we conducted collimated measurements at each depth to verify the scintillation properties across the entire crystal and the depth reconstruction. A fan beam is produced by using two lead bricks separated by an American penny with a ${}^{137}\mathrm{Cs}$ source placed on the other end. The assembly is mounted on a robotic gantry that is programmed to move in $5\ \mathrm{mm}$ steps and scan along the length of the GAGG crystal. Fig. 10 illustrates the experimental setup.

### v2

@@ Abstract
Gamma-ray astrophysics in the MeV band is an exciting field in astronomy due to its potential for multi-messenger astrophysics. It has, however, remained under-explored when compared to other wavelengths. One reason for this observational gap is the difficulties with measuring these high-energy photons and the requirement of large amounts of detection material. In this work, we investigate the usage of large-volume GAGG scintillators for use as a calorimeter in future MeV telescopes. We developed a $5\times 5$ array calorimeter utilizing $1\times 1\times 6\ \mathrm{cm}^{3}$ GAGG crystals with C-series SiPM readout. We tested the calorimeter at the High Intensity Gamma-ray Facility (HIGS) with monoenergetic beams ranging from $2-25\ \mathrm{MeV}$. Finally, we also investigated larger $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals and characterize their response across their depth when their surface treatment are either polished or frosted.
@@ 1 Introduction
The gamma-ray astrophysical sky in the MeV range remains largely under-explored. The most recent large-scale mission dedicated to measuring and imaging the MeV gamma-ray sky was the Compton Gamma-Ray Observatory (CGRO), which was de-orbited in 2000 [1]. Since then, the community has proposed several Compton and/or pair-production instruments to address what is known as the ‘MeV gap’. Most pair-production instruments are based on a two-subsystem design: a tracker to trace out the movement of the cascade of particles from the gamma-ray interaction and a calorimeter to capture the ensuing electromagnetic showers. This tracker-calorimeter design is utilized by current missions such as the Fermi Space Telescope-LAT [2] and AGILE [3]. Missions that are in the concept and development phase include AMEGO-X [4], ASTROGAM [5], SMILE-2+ [6], VLAST [7], and GECCO [8].
@@
GAGG [19] CsI:Tl[20]
Light Yield (ph/keV) 60 54
@@
Density (g/$\mathrm{cm}^{3}$) 6.6 4.51
Decay Time (ns) $\leq$150 1000
@@ 2 Development of a $5\times 5$ array GAGG Calorimeter
In this work, we utilize high light-yield GAGG from Kinheng. The prototype in this section uses $1\times 1\times 6\ \mathrm{cm}^{3}$ crystals, each one wrap with a diffuse white reflector (Tetratex) [21], and place a $6\times 6\ \mathrm{mm}^{2}$ C-series SiPM by onsemi [22] on each end of the crystal. Fig. 1 shows the bare crystal and a fully developed crystal element.
@@ 3 High-Energy Gamma-Ray Response of the Calorimeter
In April 2022, the calorimeter was taken to the Triangle Universities Nuclear Laboratory (TUNL) High Intensity Gamma-Ray Source (HIGS) facility. The facility can produce a monoenergetic gamma-ray beam with a diameter of $0.5\ \mathrm{cm}$. The experiment utilized beam energies of 2, 5, 7, 15, and 25 MeV. Fig. 5 shows the experimental setup. Since this experiment was conducted as an add-on to the AMEGO/ComPair campaign [25], the GAGG calorimeter is placed downstream of the ComPair instrument. The ComPair instrument consists of 3 subsystems: 6 layers of $0.5\ \mathrm{mm}$ of Si, $2\ \mathrm{cm}$ of CZT, and $8.35\ \mathrm{cm}$ of CsI.
@@
During the experiment, we placed the calorimeter in different positions and angles relative to the beam. This is to change the interaction profile of the gammas with the calorimeter. Fig. 6 shows the summed response to different irradiation configurations of a 5 MeV beam. Side irradiation refers to when the beam is shot into the center column of the calorimeter while front irradiation is where the beam is aimed into the $1\times 1\ \mathrm{cm^{2}}$ face of the center crystal in the array. We note that the difference in the full calorimeter response is relatively invariant in that the ratio between the escape peaks, continuum, and full-energy peak are very similar. This is likely due to the near-cubic geometry of the calorimeter which results in a uniform sensitivity to source location.
@@
Figure 7: Crystal-to-crystal response to a 5 MeV beam while the calorimeter is oriented at different poses.
Fig. 8 plots all the acquired spectra from the HIGS campaign. Full energy peaks are clearly visible for the 2, 5, and 7 $\mathrm{MeV}$ beams. Also featured are the single and double escape peaks. However, runs with 15 and 25 $\mathrm{MeV}$ do not present any evidence of escape peak, likely due to the low full-energy efficiency and full-containment of the ensuing spallation. The experimental response matches the simulated one, presented in A.
@@ 4.1 Collimated Depth Measurements
In this section, we conducted collimated measurements at each depth to verify the scintillation properties across the entire crystal and the depth reconstruction scheme. A fan beam is produced by using two lead bricks separated by an American penny with a ${}^{137}\mathrm{Cs}$ source placed on the other end. The assembly is mounted on a robotic gantry that is programmed to move in $5\ \mathrm{mm}$ steps and scan along the length of the GAGG crystal. Fig. 10 illustrates the experimental setup.